The table below shows the number of male and female students enrolled in nursing at a particular university for a recent semester. (a) Find the probability that a randomly selected student is male, given that the student is a nursing major (b) Find the probability that a randomly selected student is a nursing major, given that the student is male. Males Females Total Nursing Majors Non-nursing majors Total 1154 2396 3550 91 733 824 1063 663 2726 (a) Find the probability that a randomly selected student is male, given that the student is a nursing major The probability is (Round to three decimal places as needed.) (b) Find the probability that a randomly selected student is a nursing major, given that the student is male The probability is Round to three decimal places as needed.)
Answer #1

a) P(Male | Nursing Major) = 91/824 = 0.110

b) P(Nursing major | Male) = 91/1154 = 0.079